Use the Formula Method to answer the following problems. Show all of your work.

         DO/DA x Qty. = DG

  1.        Available are Furosemide 20 mg tablets. The order is for Furosemide 60 mg. How many tablets should be given with each dose? (2)

  1.        If a medication is available in 0.05mg tablets and the amount ordered is 0.025mg tablets, how many tablets would be given? (2)

  1.        A doctor prescribes 10 mg of cough syrup. Available is 5 mg/5 ml. How many teaspoons would you give? (3)

  1.        Zoloft 125 mg po is ordered. The strength available is Zoloft 50 mg scored tablet. How many tablets would provide the desired amount? (2)
  1.        Metformin 750 mg po is ordered. The strength available is Metformin 0.5 g tablet. How many tablet(s) would be necessary to fill the desired order? (3)
  1.        Amoxicillin oral suspension 500 mg po is ordered. Available is Amoxicillin 125 mg/5ml. a. How many ml would be administered to the client? (2)

          b. How many teaspoons would be administered? (1)

  1.        Ordered: Glucophage 1 g po with meals. Available: 500 mg tablets.

          Dose to be given: (3)

  1.        Ordered: Theophylline 300 mg po 1 h pc q 6 h. Available 200 mg tablets.

          Dose to be given: (2)

  1.        Ordered: Dipyridamole 125 mg po qid. Available 50 mg tablets.

          a. Dose to be given: (2)

           b. What is the daily dosage? (1)

  1.      Ordered: Prednisone 80 mg po daily. Available 5 mg/ml

          Dose to be given in ml: (2)

  1.      Ordered: Phenytoin 400 mg po daily. Available 100 mg capsules.

          Dose to be given: (2)

12. Ordered: Griseofulvin 1g po daily. Available: 250 mg tablets. What is the dose to be given? (3)
Answer #1

DO/DA x Qty. = DG

  1.        Available are Furosemide 20 mg tablets. The order is for Furosemide 60 mg. How many tablets should be given with each dose? (2)

One tablet is 20mg

For giving 60 mg , find how many tablets are to be given?

60/20 =3

3 tablets per dose

  1.        If a medication is available in 0.05mg tablets and the amount ordered is 0.025mg tablets, how many tablets would be given? (2)

The number of tablets to be given = dose ordered xamount in hand dose in hand

Dose ordered = 0.025mg

Dose in hand= 0.05mg

Amount in hand = 1 tablet

   tablets to be given per dose    =    0.025x1 0.05

   = 0.5

0.5 of the tablet or half the tablet has to be given.

3. A doctor prescribes 10 mg of cough syrup. Available is 5 mg/5 ml. How many teaspoons would you give? (3)

Dose ordered = 10mg

Dose in hand= 5mg

Amount in hand = 5 ml

Calculating with the same formula,

The amount of syrup to be given=    10x5 5

                                                      = 10 ml

1 teaspoon=4.9 ml or 5 ml

   Two teaspoons of the syrup has to be given

  1.        Zoloft 125 mg po is ordered. The strength available is Zoloft 50 mg scored tablet. How many tablets would provide the desired amount? (2)

Dose ordered = 125mg

Dose in hand= 50mg

Amount in hand = 1 tablet

How many tablet of Zoloft has to be given = 125x1 50  

   =2.5

Two and a half tablets has to be given per dose

  1.        Metformin 750 mg po is ordered. The strength available is Metformin 0.5 g tablet. How many tablet(s) would be necessary to fill the desired order? (3)

Dose ordered = 750mg

Dose in hand= 0.5g   =500mg

Amount in hand = 1 tablet

How many tablet has to be given = 750X1 500  

=1.5

One and half of tablet has to be given

  1.        Amoxicillin oral suspension 500 mg po is ordered. Available is Amoxicillin 125 mg/5ml. a. How many ml would be administered to the client? (2)

          b. How many teaspoons would be administered? (1)

Dose ordered = 500mg

Dose in hand= 125mg

Amount in hand = 5ml

How many ml has to be given = 500x5 125

   = 20 ml

20 ml has to be given

1teaspoon = 5ml

So   20/5 = 4 teaspoons has to be given

  1.        Ordered: Glucophage 1 g po with meals. Available: 500 mg tablets. Dose to be given: (3)

Dose ordered = 1g = 1000 mg

Dose in hand= 500mg

Amount in hand = 1 tablet

How many tablets has to be given for the dose = 1000x1 500

= 2

Two tablets has to be given

  

  1.        Ordered: Theophylline 300 mg po 1 h pc q 6 h. Available 200 mg tablets.

          Dose to be given: (2)

Dose ordered = 300 mg

Dose in hand= 200mg

Amount in hand = 1 tablet

How many tablets has to be given = 300x1 200

   =1.5

One and a half of the tablet has to be given

  1.        Ordered: Dipyridamole 125 mg po qid. Available 50 mg tablets.
    1. Dose to be given: (2)

Dose ordered = 125 mg

Dose in hand= 50mg

Amount in hand = 1 tablet

Dose to be given = 125x1 50

   = 2.5

Two and half of tablets has to be given.

  1. What is the daily dosage? (1)

The medicine has to be given 4 times a day

The daily dosage = 125×4 = 500 mg

The amount of tablets needed per day= 2.5×4 = 10 tablets

  1.      Ordered: Prednisone 80 mg po daily. Available 5 mg/ml

          Dose to be given in ml: (2)

Dose ordered = 80 mg

Dose in hand= 5mg

Amount in hand = 1 ml

Dose in ml =

80x1 5

=16 ml

16 ml has to be given per dose

  1.      Ordered: Phenytoin 400 mg po daily. Available 100 mg capsules.

          Dose to be given: (2)

Dose ordered = 400 mg

Dose in hand= 100mg

Amount in hand = 1 capsule

Number of capsules per dose = 400x1 100

   = 4

4 capsules has to be given per dose

  1. Ordered: Griseofulvin 1g po daily. Available: 250 mg tablets. What is the dose to be given? (3)

Dose ordered = 1g =1000 mg

Dose in hand= 250mg

Amount in hand = 1 tablet

How many tablet per dose has to be given = 1000x1 250 = 4

4 tablets has to be given per dose